OpenStudy (anonymous):

how do you simplify this 2 6/72

OpenStudy (anonymous):

I'm going to go forward and assume that this is: \[2 \times \frac{ 6 }{ 72 }\] At which point the 2 is multiplied times the 6 so you receive: \[\frac{ 12 }{ 72 }\] Now then, you want to simplify this fraction. A common number that both go into is 12, but you can start out with 2 and work your way up to 12 if you don't immediately see 12. 12/12=1 72/12=6 Therefore: \[\frac{ 12 }{ 72 }=\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

it is 2 wholes and 6 over 72

OpenStudy (anonymous):

So: \[2+\frac{ 6 }{ 72 }\] ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

... Well ... sorry. I don't know what is going on.

OpenStudy (anonymous):

\[2\frac{ 6}{ 72}\]

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

2+6/72=144/72+6/72=150/72

OpenStudy (anonymous):

i need help on this one instead \[1\frac{ 3}{ 8}*2\frac{ 2}{ 9}\]

OpenStudy (anonymous):

Well, 1 = x/8? In other words, how many 8's are in 1 when divided by 8? Add that number to 3. Do the same thing for 2 +2/9 and then multiply the combined top and bottom.